OpenStudy (anonymous):

PLZ HELP WITH THREE QUESTIONS. I WILL GIVE A MEDAL, I TRIED BUT I GOT THE WRONG ANSWER

OpenStudy (anonymous):

Question 6 (Multiple Choice Worth 2 points) twenty-three equals zero point five times b This equation could be used to answer which question? twenty-three is five percent of what number? What is twenty-three percent of zero point five? fifty percent of what number is twenty-three? What is fifty percent of twenty-three? Question 7 (Multiple Choice Worth 2 points) forty-five is what percent of fifteen? 3 thirty thirty-three three hundred Question 8 (Multiple Choice Worth 2 points) "What is ninety percent of two hundred fifty?" Which equation could be used to answer this question? x equals zero point nine times two hundred fifty ninety equals two hundred fifty x x equals zero point zero nine times two hundred fifty two hundred fifty equals x times ninety

OpenStudy (anonymous):

PLZ HELP, I DONT GET IT AND MY TEACHER WONT HELP ME :(

OpenStudy (anonymous):

the first one is fifty percent of what number is twenty-three?

OpenStudy (anonymous):

You need to reread the chapter on this material; I'll help, but it's obvious that you don't understand percentages.

OpenStudy (anonymous):

45 = x * 15

OpenStudy (anonymous):

second one is 300 percent

OpenStudy (anonymous):

I AM SORRY

OpenStudy (anonymous):

LOOK I CANT READ PERFECT

OpenStudy (anonymous):

45/15 = x * 15/15 3 = x

OpenStudy (anonymous):

plz dont judge

OpenStudy (anonymous):

300%

OpenStudy (anonymous):

3rd is 225

OpenStudy (anonymous):

ok and thats all?

OpenStudy (anonymous):

umm..... wait

OpenStudy (anonymous):

there is no 225 in the 3rd?

OpenStudy (anonymous):

like is it a,b,c,or d?

OpenStudy (anonymous):

look at my posts

OpenStudy (anonymous):

i am

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

cuz u only posted for the 1st and 2nd question

OpenStudy (anonymous):

x equals zero point nine times two hundred fifty

OpenStudy (anonymous):

thats for the last one

OpenStudy (anonymous):

im certain

OpenStudy (anonymous):

ok thank you so much ur the best

OpenStudy (anonymous):

medal? Silvihappy7: PLZ HELP WITH THREE QUESTIONS. I WILL GIVE A MEDAL, I TRIED BUT I GOT THE WRONG ANSWER